OpenStudy (anonymous):

Help with converting the equation to rectangular form? cos2Ө=1

hartnn (hartnn):

whats your progress ? did you try ?

OpenStudy (anonymous):

cos2Ө=1 (cos^2Ө)-(sin^2Ө)=1 (using the double angle formula) But from here, I'm stuck :(

hartnn (hartnn):

basically you'll need these : \(\cos 2\theta = \cos^2 \theta- \sin^2 \theta\) then divide both sides by cos^2 theta

hartnn (hartnn):

then use \(\sec^2\theta = 1+ \tan^2 \theta\)

hartnn (hartnn):

and lastly, \(\theta = \tan^{-1}\dfrac{y}{x}\)

hartnn (hartnn):

tell us what u get , so that we can verify your steps...

OpenStudy (anonymous):

cos2Ө=1 (cos^2Ө)-(sin^2Ө)=1 1-(tan^2Ө)=(1/cos^2Ө).....divided both sides by (cos^2Ө) 1-(tan^2Ө)=(sec^2Ө) Is this correct so far? ... sorry I took so long to reply,, I appreciate you helping me

hartnn (hartnn):

yes, correct, so far...its ok. now, (sec^2Ө) = 1+ (tan^2Ө)

OpenStudy (anonymous):

I think I made a mistake because I got 1-(tan^2Ө)=(sec^2Ө) instead of "+" like you wrote

hartnn (hartnn):

no, you did not make any mistake. 1-(tan^2Ө)=(sec^2Ө) then since, (sec^2Ө) = 1+ (tan^2Ө) we get 1-(tan^2Ө) = 1+ (tan^2Ө) isolate (tan^2Ө) from here, and put theta=...

OpenStudy (anonymous):

1-(tan^2Ө)=(sec^Ө) 1-(tan^2Ө)=1+(tan^2Ө) 0=2(tan^2Ө) 0=(tan^2Ө) 0=tanӨ 0=y/x y=0(x) y=0....final answer

OpenStudy (anonymous):

Thank you for taking time to help me,, I appreciate it!

hartnn (hartnn):

y=0 is correct, but don't forget to mention that \(x \ne 0\)